OpenStudy (anonymous):

i have problems with word problems....a girl drove 250 miles in the same amount of time it took a turbopeller plane to travel 1050 miles. The speed of the plane was 160 mph faster than the speed of the car. find the speed of the plane?? HELP

OpenStudy (anonymous):

ok, you know s=v*t where s is distance, v is speed, t is time. lets draw a table car plane s 250 1050 v x x+160 t t = t t=s/v so 250/x = 1050/(x+160) then solve this equation :)

OpenStudy (anonymous):

your a lifesaver! thanks!

OpenStudy (anonymous):

np :D